In a large city, two newspapers are published, the Sun and the Post. The circulation departments report than 22% of the city's households have a subscription to the Sun and 35% subscribe to the Post.

1. What is the probability of a household subscribing to both the Sun and the Post?

2. What is the probability of a household subscribing to either the Sun or the Post?
Answer #1

1) considering independence:

probability of a household subscribing to both the Sun and the Post

=P(subs to Sun)*P(subs to Post)=0.22*0.35=0.0770

2)

probability of a household subscribing to either the Sun or the Post

=P(Sun)+P(Post)-P(both)=0.22+0.35-0.0770=0.4930